OpenStudy (anonymous):

Explain the connection between angles and complex numbers we've been relying on. Illustrate this by finding the sine and cosine of 1. A 22.5 degree angle, and 2. twice the angle in a 5-12-13 triangle opposite the side of length 5.

OpenStudy (amistre64):

if we rename the axises of a conventional graph: Y to i, and X to Real; we have a representation of the complex plane. the complex number a+bi can then be positioned on the plane as (r,i) coordinants: r=a, i=b any point on a plane can be defined by its distance and movement from the origin

OpenStudy (anonymous):

ya i think i have this, can i ask you something though? for twice the angle, is it okay to use de moivre's theorem?

OpenStudy (anonymous):

it sounds right but looks wrong i'm not quite sure...

OpenStudy (amistre64):

not sure what DM looks like in this, but my interpretation of the question is something like this 

OpenStudy (amistre64):

given an abc right triangle \[sin\frac{\alpha}{2}=\frac{c-b}{\sqrt{2c(c-b)}}\] \[cos\frac{\alpha}{2}=\frac{a}{\sqrt{2c(c-b)}}\] but i might have those mixed up

OpenStudy (amistre64):